Sale of Asset

Equipment acquired on January 9, 20Y3, at a cost of $731,000, has an estimated useful life of 16 years, an estimated residual value of $87,720, and is depreciated by the straight-line method.

a. What was the book value of the equipment at the end of the fifth year, December 31, 20Y7? Round your interim calculations and final answer to the nearest dollar.
$

For decreases in accounts or outflows of cash, enter your answers as negative numbers. Round annual depreciation to the nearest dollar and use this amount in your follow-on calculations. If no account or activity is affected, select "No effect" from the dropdown and leave the corresponding number entry box blank.

b1. Assuming that the equipment was sold on July 1, 20Y8, for $292,400, illustrate the effects on the accounts and financial statement of depreciation for the six months until the sale date.
Answer #1

Solution

Sale of asset –

  1. Determination of book value of the equipment at the end of fifth year, December 31, 20Y7:

Cost of equipment = $731,000

Residual value = $87,720

Depreciable base = $643,280

Useful life = 16 years

Depreciation expense under straight line method = depreciable base x 1/useful life

Annual depreciation expense = $643,280 x 1/16 = $40,205

Since annual depreciation expense under straight line method remains same throughout the useful life of the asset, the depreciation expense for each year is $40,205.

Book value = cost – accumulated depreciation

Book value at the end of fifth year = cost - accumulated depreciation for five years

Accumulated depreciation for five years = 40,205 x 5 = $201,025

Book value = 731,000 – 201,025 = $529,975

Hence, book value of the equipment at the end of the fifth year, December 31, 20Y7 = $529,975

  1. Demonstrating the effects of sale of equipment on July 1, 20Y8 for $292,400 –

Depreciation expense for the six months, Jan 1, 20Y8 – July 1, 20Y8 = $40,205 x 6/12 = $20,103

Book value the date of sale = cost – accumulated depreciation

Accumulated depreciation = depreciation up to five years + depreciation for 6 month in 20Y8

= $201,025 + $20,103 = $221,128

Gain on sale of equipment = $292,400 - $221,128 = $71,272

When the asset is sold,

The income increases by gain of $71, 272

The cash flow increases by $292,400

Date

Account Titles and Explanation

Debit

Credit

July 1, 20Y8

Depreciation Expense

$20,103

Accumulated Depreciation

$20,103

(To record depreciation expense for 6 months)

July 1, 20Y8

Cash

$292,400

Accumulated Depreciation

$221,128

Gain on sale of equipment

$71,272

(To record sale of equipment and gain on sale)